The following table contains information concerning four jobs that are awaiting processing at a work center.

Job Job Time (Days) Due Date
(Days)
A 14 20
B 10 16
C 7 15
D 6 17


a. Sequence the jobs using (1) First come, first served, (2) Shortest processing time, (3) Earliest due date, and (4) Critical ratio. Assume the list is by order of arrival.

Sequence for First come, first served   (Click to select)   A-B-C-D   D-C-B-A   C-B-D-A   A-C-D-B
Sequence for Shortest processing time   (Click to select)   A-C-D-B   C-B-D-A   D-C-B-A   A-B-C-D
Sequence for Earliest due date   (Click to select)   A-C-D-B   D-C-B-A   A-B-C-D   C-B-D-A
Sequence for Critical ratio   (Click to select)   A-C-D-B   A-B-C-D   D-C-B-A   C-B-D-A


b. For each of the methods in part a, determine (1) the average job flow time, (2) the average tardiness, and (3) the average number of jobs at the work center. (Round your intermediate calculations and final answers to 2 decimal places.)
  

First Come, First
Served
Shortest Processing
Time
Earliest Due Date Critical Ratio
Average job flow time
Average tardiness
Average number of jobs
Answer #1

a)

Sequence for FCFS = A-B-C-D

Sequence for SPT = D-C-B-A
Sequence for EDD= C-B-D-A
Sequence for Critical ratio = A-B-C-D

b.

Formula